# Get rich quick: exact solutions reveal how unbalanced initializations promote rapid feature learning

### v1

@@ 4 Wide and Deep Linear Networks
We now show how the analysis techniques used to study the influence of initialization geometry in the single-neuron setting can be applied to linear networks with multiple neurons, outputs, and layers.
Wide linear networks. We consider the dynamics of a two-layer linear network with $h$ hidden neurons and $c$ outputs, $f(x;\theta)=A^{\intercal}Wx$, where $W\in\mathbb{R}^{h\times d}$ and $A\in\mathbb{R}^{h\times c}$. We assume that $\min\{d,c\}\leq h\leq\max\{d,c\}$, such that this parameterization can represent all linear maps from $\mathbb{R}^{d}\to\mathbb{R}^{c}$. As in the single-neuron setting, the rescaling symmetry in this model implies that $\Delta=\eta_{w}A_{0}A_{0}^{\intercal}-\eta_{a}W_{0}W_{0}^{\intercal}$ remains conserved throughout gradient flow [62]. The NTK matrix is $K=\left(\mathbf{I}_{c}\otimes X\right)\left(\eta_{w}A^{\intercal}A\oplus\eta_{a}W^{\intercal}W\right)\left(\mathbf{I}_{c}\otimes X^{\intercal}\right)$, where $\otimes$ and $\oplus$ denote the Kronecker product and sum [^4] The Kronecker sum is defined for square matrices $A\in\mathbb{R}^{c\times c}$ and $B\in\mathbb{R}^{d\times d}$ as $A\oplus B=A\otimes\mathbf{I}_{d}+\mathbf{I}_{c}\otimes B$. respectively. Drawing insights from our analysis of the single-neuron scenario ($h=c=1$), we consider the dynamics of $\beta=W^{\intercal}A\in\mathbb{R}^{d\times c}$,
@@
where $\mathrm{vec}(\cdot)$ denotes the vectorization operator. As in the single-neuron setting, we find that the dynamics of $\beta$ are preconditioned by a matrix $M$ that depends on quadratics of $A$ and $W$. $M$ can be expressed in terms of the matrices $\beta_{i}=w_{i}a_{i}^{\intercal}\in\mathbb{R}^{d\times c}$, which represent the contribution to the input-output map of a single hidden neuron with parameters $w_{i}$, $a_{i}$ and conserved quantity $\delta_{i}=\Delta_{ii}$.
@@
Assume $\|\beta_{i}\|_{F}\neq 0$ for all $i\in[h]$ and let $\kappa_{i}=\sqrt{\delta_{i}^{2}+4\eta_{a}\eta_{w}\|\beta_{i}\|_{F}^{2}}$, then the matrix $M$ can be expressed as the sum $M=\sum_{i=1}^{h}M_{i}$ over hidden neurons where $M_{i}$ is defined as,
@@
$M_{i}=\left(\frac{\kappa_{i}+\delta_{i}}{2}\right)\frac{\beta_{i}^{\intercal}\beta_{i}}{\|\beta_{i}\|_{F}^{2}}\oplus\left(\frac{\kappa_{i}-\delta_{i}}{2}\right)\frac{\beta_{i}\beta_{i}^{\intercal}}{\|\beta_{i}\|_{F}^{2}}.$ (7)
By studying the dependence of $M$ on the conserved quantities $\delta_{i}$ and the relative sizes of dimensions $d$, $h$ and $c$, we can determine the influence of the initialization geometry on the learning regime [^5] As shown in Appendix B, we can recover Eq. 4 presented in the single-neuron setting directly from Eq. 7..
@@
Funnel and inverted-funnel networks. We consider funnel networks, which narrow from input to output ($d>h\geq c$), and inverted-funnel networks, which expand from input to output ($d\leq h<c$). Except for a measure zero set of initializations, funnel networks always enter the lazy regime as all $\delta_{i}\rightarrow\infty$, whereas inverted funnel networks do so as $\delta_{i}\to-\infty$. As elaborated in Appendix B, this occurs because in these limits a solution $\beta_{*}$ must exist within the space spanned by $M$ at initialization. In the opposite limits of $\delta$, these networks enter a lazy followed by rich regime, assuming $\beta_{0}\not\propto\beta_{*}$. Conversely, as the $\delta_{i}\to 0$, all networks transition into the rich regime.
[…]
@@
Deep linear networks. As detailed in Section B.2, we generalize the inductive bias derived for rich two-layer linear networks by Azulay et al. [9] to encompass deep linear networks. For a depth-$(l+1)$ linear network, $f(x;\theta)=A^{\intercal}\prod_{i=1}^{l}W_{i}x$, where $\beta=\prod_{i=1}^{l}W_{i}^{\intercal}A$, we find that the inductive bias of the rich regime is $Q(\beta)=(\tfrac{l+1}{l+2})\|\beta\|^{\frac{l+2}{l+1}}-\|\beta_{0}\|^{-\frac{l}{l+1}}\beta_{0}^{\intercal}\beta$ (see Theorem B.9). This inductive bias strikes a balance between attaining the minimum norm solution and preserving the initialization direction, which with increased depth emphasizes the latter.
@@ 5 Piecewise Linear Networks
We now take a first step towards extending our analysis from linear networks to piecewise linear networks with activation functions of the form $\sigma(z)=\max\{z,\gamma z\}$. The input-output map of a piecewise linear network with $l$ hidden layers and $h$ hidden neurons per layer is comprised of potentially $O(h^{dl})$ convex activation regions [63]. Each region is defined by a unique activation pattern of the hidden neurons. The input-output map is linear within each region and continuous at the boundary between regions. Collectively, the activation regions form a 2-colorable [^6] To our knowledge, this property has not been recognized before. See Section C.1 for a formal statement. convex partition of input space, as shown in Fig. 5. We investigate how the initialization geometry influences the evolution of this partition and the linear maps within each region.
Figure 5: Evolution of a ReLU network. Here we depict the surface of a two-layer ReLU network with twenty hidden neurons, starting from a nearly zero input-output map $f(x;\theta_{0})\approx 0$, trained on a task akin to exclusive or. The surface consists of convex conic regions, each with a distinct activation pattern, colored by the parity of active neurons. With a lazy downstream initialization (bottom), the activation partition is fixed, and the network fits the data by reweighing the hidden neurons. With a rich balanced or upstream initialization (top), the network undergoes an initial alignment phase where the partition map changes rapidly while the input-output map remains close to zero, followed by a data fitting phase.
@@
Two-layer network. We consider the dynamics of a two-layer piecewise linear network without biases, $f(x;\theta)=a^{\intercal}\sigma(Wx)$, where $W\in\mathbb{R}^{h\times d}$ and $a\in\mathbb{R}^{h}$. Following the approach in Section 4, we consider the contribution to the input-output map from a single hidden neuron $i\in[h]$ with parameters $w_{i}\in\mathbb{R}^{d}$, $a_{i}\in\mathbb{R}$ and conserved quantity $\delta_{i}=\eta_{w}a_{i}^{2}-\eta_{a}\|w_{i}\|^{2}$ [62]. However, unlike the linear setting, the neuron’s contribution to $f(x_{j};\theta)$ is regulated by whether the input $x_{j}$ is in the neuron’s active halfspace. Let $C\in\mathbb{R}^{h\times n}$ be the matrix with elements $c_{ij}=\sigma^{\prime}(w_{i}^{\intercal}x_{j})$, which determines the activation of the $i^{\mathrm{th}}$ neuron for the $j^{\mathrm{th}}$ data point, and let $\rho_{j}=f(x_{j};\theta)-y_{j}$. Then we can express the dynamics of $\beta_{i}=a_{i}w_{i}$ as,
@@
$\dot{\beta}_{i}=-\underbrace{\left(\eta_{w}a_{i}^{2}I_{d}+\eta_{a}w_{i}w_{i}^{\intercal}\right)}_{M_{i}}\underbrace{\textstyle\sum_{j=1}^{n}c_{ij}x_{j}\rho_{j}}_{\xi_{i}}.$ (8)
@@
The matrix $M_{i}\in\mathbb{R}^{d\times d}$ is a preconditioning matrix on the dynamics, and when $\beta_{i}\neq 0$, it can be expressed in terms of $\beta_{i}$ and $\delta_{i}$. Unlike the linear setting, $\xi_{i}\in\mathbb{R}^{d}$ driving the dynamics is not shared for all neurons because of its dependence on $c_{ij}$. Additionally, the NTK matrix in this setting depends on $M_{i}$ and $C$, with elements $K_{jk}=\sum_{i=1}^{h}c_{ij}x_{j}^{\intercal}M_{i}x_{k}c_{ik}$. To examine the evolution of $K$, we consider a signed spherical coordinate transformation separating the dynamics of $\beta_{i}$ into its directional $\hat{\beta}_{i}=\mathrm{sgn}(a_{i})\tfrac{\beta_{i}}{\|\beta_{i}\|}$ and radial $\mu_{i}=\mathrm{sgn}(a_{i})\|\beta_{i}\|$ components, such that $\beta_{i}=\mu_{i}\hat{\beta}_{i}$. $\hat{\beta}_{i}$ determines the direction and orientation of the halfspace where the $i^{\mathrm{th}}$ neuron is active, while $\mu_{i}$ determines the slope of the contribution in this halfspace. These coordinates evolve according to,
@@
$\dot{\mu}_{i}=-\sqrt{\delta_{i}^{2}+4\eta_{a}\eta_{w}\mu_{i}^{2}}\hat{\beta}_{i}^{\intercal}\xi_{i},\qquad\dot{\hat{\beta}}_{i}=-\frac{\sqrt{\delta_{i}^{2}+4\eta_{a}\eta_{w}\mu_{i}^{2}}+\delta_{i}}{2\mu_{i}}\left(I_{d}-\hat{\beta}_{i}\hat{\beta}_{i}^{\intercal}\right)\xi_{i}.$ (9)
@@
Downstream. When $\delta_{i}\ll 0$, $M_{i}\approx|\delta_{i}|\hat{\beta}_{i}\hat{\beta}_{i}^{\intercal}$. The dynamics are approximately $\partial_{t}\hat{\beta}_{i}=0$ and $\partial_{t}\mu_{i}=-|\delta_{i}|\hat{\beta}_{i}^{\intercal}\xi_{i}$. Irrespective of $\xi_{i}$, $\hat{\beta}_{i}(t)=\hat{\beta}_{i}(0)$, which implies the overall partition map doesn’t change (Fig. 5, bottom), nor the activation patterns $C$, nor $M_{i}$. Only $\mu_{i}$ changes to fit the data, while the NTK remains constant. If the number of hidden neurons is insufficient to fit the data, there will be a second, slow rich alignment phase where the kernel will change, with $|\delta_{i}|$ controlling the delay.
@@
Balanced. When $\delta_{i}=0$, $M_{i}=\sqrt{\eta_{a}\eta_{w}}|\mu_{i}|(\mathbf{I}_{d}+\hat{\beta}_{i}\hat{\beta}_{i}^{\intercal})$, and the dynamics simplify to, $\partial_{t}\hat{\beta}_{i}=-\sqrt{\eta_{a}\eta_{w}}\mathrm{sgn}(\mu_{i})(\mathbf{I}_{d}-\hat{\beta}_{i}\hat{\beta}_{i}^{\intercal})\xi_{i}$ and $\partial_{t}\mu_{i}=-2\sqrt{\eta_{a}\eta_{w}}|\mu_{i}|\hat{\beta}_{i}^{\intercal}\xi_{i}$. Here both the direction and magnitude of $\beta_{i}$ evolve, resulting in changes to the activation regions, patterns $C$, and NTK $K$. For vanishing initializations where $\|\beta_{i}(0)\|\to 0$ for all $i\in[h]$, we can decouple the dynamics into two distinct phases of training, analogous to the rich regime discussed in Section 3 and shown in Fig. 5, top. Phase I: Partition alignment. At vanishing scale, the output $f(x;\theta_{0})\approx 0$ for all input $x$, such that the vector driving the dynamics $\xi_{i}\approx-\sum_{j=1}^{n}c_{ij}x_{j}y_{j}$ is independent of the other hidden neurons. At the same time, the radial dynamics slow down relative to the directional dynamics, and the function’s output will remain small as each neuron aligns decoupled from the rest. Prior works have introduced structural constraints on the training data, such as orthogonally separable [50, 53, 54], pair-wise orthonormal [52], linearly separable and symmetric [51] or small angle [55], to analytically determine the fixed points of this alignment phase. Phase II: Data fitting. After enough time, the magnitudes of $\beta_{i}$ have grown such that we can no longer assume $f(x;\theta)\approx 0$ and thus the residual will depend on all $\beta_{i}$. In this phase, the radial dynamics dominate the learning driving the network to fit the data. However, it is possible for the directions to continue to change, and thus prior works have further decomposed this phase into multiple stages.
@@
Upstream. When $\delta_{i}\gg 0$, $M_{i}\approx\delta_{i}\mathbf{I}_{d}$, and the dynamics are approximately $\partial_{t}\hat{\beta}_{i}=-\delta_{i}\mu_{i}^{-1}(\mathbf{I}_{d}-\hat{\beta}_{i}\hat{\beta}_{i}^{\intercal})\xi_{i}$ and $\partial_{t}\mu_{i}=-\delta_{i}\hat{\beta}_{i}^{\intercal}\xi_{i}$. Again, both the direction and magnitude of $\beta_{i}$ change. However, unlike the balanced setting, in this setting $M_{i}$ is independent of $\beta_{i}$ and stays constant through training. Yet, as $\beta_{i}$ change in direction, so can $C$, and thus the NTK. This setting is unique, because it is rich due to a changing activation pattern, but the dynamics do not move far in parameter space. Furthermore, unlike in the balanced scenario where scale adjusts the speed of radial dynamics, here it regulates the speed of directional dynamics, with vanishing initializations prompting an extremely fast alignment phase, as observed in Fig. 1.
@@
Figure 6: Impact of upstream initializations in practice. Here we provide evidence that an upstream initialization (a) drives feature learning through changing activation patterns, (b) promotes interpretability of early layers in CNNs, (c) reduces the sample complexity of learning hierarchical data, and (d) decreases the time to grokking in modular arithmetic. In these experiments, we regulate the learning speed of the first layer relative to the rest of the network by scaling the initialization without altering the input-output map. Here, $\alpha=1$ denotes a standard parameterization. $\alpha\gg 1$ and $\alpha\ll 1$ correspond to upstream and downstream initializations, respectively. See Section D.3.

### v2

@@ 4 Wide and Deep Linear Networks
We now show how the analysis techniques used to study the influence of relative scale in the single-neuron setting can be applied to linear networks with multiple neurons, outputs, and layers.
Wide linear networks. We consider the dynamics of a two-layer linear network with $h$ hidden neurons and $c$ outputs, $f(x;\theta)=A^{\intercal}Wx$, where $W\in\mathbb{R}^{h\times d}$ and $A\in\mathbb{R}^{h\times c}$. We assume $h\geq\min(d,c)$, such that this parameterization can represent all linear maps from $\mathbb{R}^{d}\to\mathbb{R}^{c}$. The rescaling symmetry between $A$ and $W$ implies the $h\times h$ matrix $\Delta=\eta_{w}A_{0}A_{0}^{\intercal}-\eta_{a}W_{0}W_{0}^{\intercal}$ is conserved throughout gradient flow [62]. Drawing insights from our analysis of the single-neuron scenario ($h=c=1$), we consider the dynamics of $\beta=W^{\intercal}A\in\mathbb{R}^{d\times c}$,
@@
where $\mathrm{vec}(\cdot)$ denotes the vectorization operator and $\oplus$ denotes the Kronecker sum [^5] The Kronecker sum is defined for square matrices $C\in\mathbb{R}^{c\times c}$ and $D\in\mathbb{R}^{d\times d}$ as $C\oplus D=C\otimes\mathbf{I}_{d}+\mathbf{I}_{c}\otimes D$.. As in the single-neuron setting, we find that the dynamics of $\beta$ are preconditioned by a matrix $M$ that depends on quadratics of $A$ and $W$ and characterizes the NTK matrix $K=\left(\mathbf{I}_{c}\otimes X\right)M\left(\mathbf{I}_{c}\otimes X^{\intercal}\right)$. We now show how $M$ can be expressed [^6] When $h=c=1$ we can recover Eq. 4 presented in the single-neuron setting directly from Eq. 7. in terms of the rank-1 matrices $\beta_{k}=w_{k}a_{k}^{\intercal}\in\mathbb{R}^{d\times c}$, which represent the contribution to $\beta$ of a single neuron with parameters $w_{k}$, $a_{k}$ and conserved quantity $\delta_{k}=\Delta_{kk}$.
@@
Whenever $\|\beta_{k}\|_{F}\neq 0$ for all $k\in[h]$, the matrix $M$ can be expressed as the sum $M=\sum_{k=1}^{h}M_{k}$ over hidden neurons where $M_{k}$ is defined as,
@@
$M_{k}=\left(\frac{\sqrt{\delta_{k}^{2}+4\eta_{a}\eta_{w}\|\beta_{k}\|_{F}^{2}}+\delta_{k}}{2}\right)\frac{\beta_{k}^{\intercal}\beta_{k}}{\|\beta_{k}\|_{F}^{2}}\oplus\left(\frac{\sqrt{\delta_{k}^{2}+4\eta_{a}\eta_{w}\|\beta_{k}\|_{F}^{2}}-\delta_{k}}{2}\right)\frac{\beta_{k}\beta_{k}^{\intercal}}{\|\beta_{k}\|_{F}^{2}}.$ (7)
@@
By studying the dependence of $M$ on the conserved quantities $\delta_{k}$ and the dimensions $d$, $h$ and $c$, we can determine the influence of the relative scale on the learning regime. When $\min(d,c)\leq h<\max(d,c)$, and assuming independent initializations for all $\beta_{k}$, then networks which narrow from input to output ($d>c$) enter the lazy regime when all $\delta_{k}\gg 0$, whereas networks which expand from input to output ($d<c$) do so when all $\delta_{k}\ll 0$. However, with opposite signs for $\delta_{k}$, and assuming all $\beta_{k}(0)\not\propto\beta_{*}$, these networks enter a delayed rich regime. As elaborated in Section B.1.5, this occurs because in these regimes a solution $\beta_{*}$ does not exist within the space spanned by $M$ at initialization. When $h\geq\max(d,c)$ all networks enter the lazy regime when all $\delta_{k}\gg 0$ or all $\delta_{k}\ll 0$. Conversely, as all $\delta_{k}\to 0$, all networks transition into the rich regime regardless of dimensions. While Theorem 4.1 offers valuable insight into the learning regimes in the limits of $\delta_{k}$, understanding the transition between regimes remains challenging. To achieve this, we aim to express $M$ in terms of $\beta$, rather than $\beta_{k}$, by introducing structure on the conserved quantities $\Delta$.
Theorem 4.2.
When $\Delta=\delta\mathbf{I}_{h}$ and $h=d$ if $\delta<0$ or $h=c$ if $\delta>0$, then the matrix $M$ can be expressed as $M=\sqrt{\eta_{a}\eta_{w}\beta^{\intercal}\beta+\frac{\delta^{2}}{4}\mathbf{I}_{c}}\otimes\mathbf{I}_{d}+\mathbf{I}_{c}\otimes\sqrt{\eta_{a}\eta_{w}\beta\beta^{\intercal}+\frac{\delta^{2}}{4}\mathbf{I}_{d}}$.
From Theorem 4.2 the resulting dynamics of $\beta$ simplify to a self-consistent equation regulated by $\delta$,
$\dot{\beta}=-X^{\intercal}P\sqrt{\eta_{a}\eta_{w}\beta^{\intercal}\beta+\frac{\delta^{2}}{4}\mathbf{I}_{c}}-\sqrt{\eta_{a}\eta_{w}\beta\beta^{\intercal}+\frac{\delta^{2}}{4}\mathbf{I}_{d}}X^{\intercal}P,$ (8)
where $P=X\beta-Y$ is the residual. Under our isotropic assumption on the conserved quanitities $\Delta=\delta\mathbf{I}_{h}$, these dynamics are exact. Concurrent to our work, Tu et al. [63] finds that $\beta$ approximately follows these dynamics in the overparameterized setting $h\gg\max(d,c)$ under a Gaussian initialization $\mathcal{N}(0,\sigma^{2})$ of the parameters where $\sigma^{2}h$ is analogous to $\delta$.
Equipped with a self-consistent equation for the dynamics of $\beta$ we now aim to interpret these dynamics as a mirror flow with a $\delta$-dependent potential. As presented in Theorem B.6, the dynamics of the singular values of $\beta$ can be described as a mirror flow with a hyperbolic entropy potential, which smoothly interpolates between an $\ell^{1}$ and $\ell^{2}$ penalty on the singular values for the rich ($\delta\to 0$) and lazy ($\delta\to\pm\infty$) regimes respectively. This potential was first identified as the inductive bias for diagonal linear networks by Woodworth et al. [14] and the same mirror flow on the singular values is derived from a different initialization choice in prior work by Varre et al. [64].
@@
Deep linear networks. As presented in Theorem B.10, we generalize the inductive bias derived for rich two-layer linear networks by Azulay et al. [9] to deep linear networks. For a depth-$(L+1)$ linear network, $f(x;\theta)=A^{\intercal}\prod_{l=1}^{L}W_{l}x$, where $\beta=\prod_{l=1}^{L}W_{l}^{\intercal}A$, we find that the inductive bias of the rich regime is $Q(\beta)=(\tfrac{L+1}{L+2})\|\beta\|^{\frac{L+2}{L+1}}-\|\beta_{0}\|^{-\frac{L}{L+1}}\beta_{0}^{\intercal}\beta$. This inductive bias strikes a depth-dependent balance between attaining the minimum norm solution and preserving the initialization direction.
@@ 5 Piecewise Linear Networks
We now take a first step towards extending our analysis from linear networks to piecewise linear networks with activation functions of the form $\sigma(z)=\max(z,\gamma z)$. The input-output map of a piecewise linear network with $L$ hidden layers and $h$ hidden neurons per layer is comprised of potentially $O(h^{dL})$ convex activation regions [65]. Each region is defined by a unique activation pattern of the hidden neurons. The input-output map is linear within each region and continuous at the boundary between regions. Collectively, the activation regions form a 2-colorable [^7] To our knowledge, this property has not been recognized before. See Section C.1 for a formal statement. convex partition of input space, as shown in Fig. 5. We investigate how the relative scale influences the evolution of this partition and the linear maps within each region.
@@
Two-layer network. We consider the dynamics of a two-layer piecewise linear network without biases, $f(x;\theta)=a^{\intercal}\sigma(Wx)$, where $W\in\mathbb{R}^{h\times d}$ and $a\in\mathbb{R}^{h}$. Following the approach in Section 4, we consider the contribution to the input-output map from a single hidden neuron $k\in[h]$ with parameters $w_{k}\in\mathbb{R}^{d}$, $a_{k}\in\mathbb{R}$ and conserved quantity $\delta_{k}=\eta_{w}a_{k}^{2}-\eta_{a}\|w_{k}\|^{2}$ [62]. However, unlike the linear setting, the neuron’s contribution to $f(x_{i};\theta)$ is regulated by whether the input $x_{i}$ is in the neuron’s active halfspace. Let $C\in\mathbb{R}^{h\times n}$ be the matrix with elements $c_{ki}=\sigma^{\prime}(w_{k}^{\intercal}x_{i})$, which determines the activation of the $k^{\mathrm{th}}$ neuron for the $i^{\mathrm{th}}$ data point. The dynamics of $\beta_{k}=a_{k}w_{k}$ are,
@@
$\dot{\beta}_{k}=-\underbrace{\left(\eta_{w}a_{k}^{2}\mathbf{I}_{d}+\eta_{a}w_{k}w_{k}^{\intercal}\right)}_{M_{k}}\underbrace{\textstyle\sum_{i=1}^{n}c_{ki}x_{i}(f(x_{i};\theta)-y_{i})}_{\xi_{k}}.$ (9)
@@
The matrix $M_{k}\in\mathbb{R}^{d\times d}$ is a preconditioning matrix on the dynamics, and when $\beta_{k}\neq 0$, it can be expressed in terms of $\beta_{k}$ and $\delta_{k}$. Unlike the linear setting, $\xi_{k}\in\mathbb{R}^{d}$ driving the dynamics is not shared for all neurons because of its dependence on $c_{ki}$. Additionally, the NTK matrix in this setting depends on $M_{k}$ and $C$, with elements $K_{ij}=\sum_{k=1}^{h}c_{ki}x_{i}^{\intercal}M_{k}x_{j}c_{kj}$. To examine the evolution of $K$, we consider a signed spherical coordinate transformation separating the dynamics of $\beta_{k}$ into its directional $\hat{\beta}_{k}=\mathrm{sgn}(a_{k})\tfrac{\beta_{k}}{\|\beta_{k}\|}$ and radial $\mu_{k}=\mathrm{sgn}(a_{k})\|\beta_{k}\|$ components, such that $\beta_{k}=\mu_{k}\hat{\beta}_{k}$. $\hat{\beta}_{k}$ determines the direction and orientation of the halfspace where the $k^{\mathrm{th}}$ neuron is active, while $\mu_{k}$ determines the slope of the contribution in this halfspace. These coordinates evolve according to,
@@
$\dot{\mu}_{k}=-\sqrt{\delta_{k}^{2}+4\eta_{a}\eta_{w}\mu_{k}^{2}}\hat{\beta}_{k}^{\intercal}\xi_{k},\qquad\dot{\hat{\beta}}_{k}=-\frac{\sqrt{\delta_{k}^{2}+4\eta_{a}\eta_{w}\mu_{k}^{2}}+\delta_{k}}{2\mu_{k}}\left(\mathbf{I}_{d}-\hat{\beta}_{k}\hat{\beta}_{k}^{\intercal}\right)\xi_{k}.$ (10)
@@
Downstream. When $\delta_{k}\ll 0$, $M_{k}\approx|\delta_{k}|\hat{\beta}_{k}\hat{\beta}_{k}^{\intercal}$, and the dynamics are approximately $\partial_{t}\hat{\beta}_{k}=0$ and $\partial_{t}\mu_{k}=-|\delta_{k}|\hat{\beta}_{k}^{\intercal}\xi_{k}$. Irrespective of $\xi_{k}$, $\hat{\beta}_{k}(t)=\hat{\beta}_{k}(0)$, which implies the overall partition map doesn’t change (Fig. 5, bottom), nor the activation patterns $C$, nor $M_{k}$. Only $\mu_{k}$ changes to fit the data, while the NTK remains constant. If the number of hidden neurons is insufficient to fit the data, there is a delayed rich alignment phase where the kernel will change, with $|\delta_{k}|$ determining the delay.
@@
Balanced. When $\delta_{k}=0$, $M_{k}=\sqrt{\eta_{a}\eta_{w}}|\mu_{k}|(\mathbf{I}_{d}+\hat{\beta}_{k}\hat{\beta}_{k}^{\intercal})$, and the dynamics simplify to, $\partial_{t}\hat{\beta}_{k}=-\sqrt{\eta_{a}\eta_{w}}\mathrm{sgn}(\mu_{k})(\mathbf{I}_{d}-\hat{\beta}_{k}\hat{\beta}_{k}^{\intercal})\xi_{k}$ and $\partial_{t}\mu_{k}=-2\sqrt{\eta_{a}\eta_{w}}|\mu_{k}|\hat{\beta}_{k}^{\intercal}\xi_{k}$. Here both the direction and magnitude of $\beta_{k}$ evolve, resulting in changes to the activation regions, patterns $C$, and NTK $K$. For vanishing initializations where $\|\beta_{k}(0)\|\to 0$ for all $k\in[h]$, we can decouple the dynamics into two distinct phases of training (Fig. 5, top), analogous to the rich regime discussed in Section 3. Phase I: Partition alignment. At vanishing scale, the output $f(x;\theta_{0})\approx 0$ for all input $x$, such that the vector driving the dynamics $\xi_{k}\approx-\sum_{i=1}^{n}c_{ki}x_{i}y_{i}$ is independent of the other hidden neurons. At the same time, the radial dynamics slow down relative to the directional dynamics, and the function’s output will remain small as each neuron aligns to certain data-dependent fixed points, decoupled from the rest. Prior works have introduced structural constraints on the training data, such as orthogonally separable [50, 53, 54], pair-wise orthonormal [52], linearly separable and symmetric [51] or small angle [55], to analytically determine the fixed points of this alignment phase. Phase II: Data fitting. After enough time, the magnitudes of $\beta_{k}$ have grown such that we can no longer assume $f(x;\theta)\approx 0$ and thus the residual will depend on all $\beta_{k}$. In this phase, the radial dynamics dominate the learning driving the network to fit the data. However, it is possible for the directions to continue to change, and therefore some prior works have further decomposed this phase into multiple stages.
(a) Evolution of a ReLU network’s input-output map
(b) Hamming and parameter distance over $\tau$-$\delta$ sweep
Figure 5: Rapid feature learning is caused by large activation changes with minimal parameter movement. (a) We show the surface of a two-layer ReLU network trained on an XOR-like task, starting with a near-zero input-output map, $f(x;\theta_{0})\approx 0$. The surface consists of convex conic regions, each with a distinct activation pattern, colored by the parity of active neurons. A lazy initialization (bottom) maintains a fixed activation partition throughout training, reweighting the hidden neurons to fit the data. In contrast, a rich balanced or upstream initialization (top) features an initial alignment phase where the partition map changes rapidly while the input-output map remains close to zero, followed by a data-fitting phase. (b) We show the evolution of Hamming distance in activation patterns and parameter distance, relative to $t=0$, as a function of overall and relative scales (same experiments as in Fig. 1(b)). Rapid feature learning occurs from a small-$\tau$ upstream initialization that promotes faster learning in early layers, driving a large change in Hamming distance, but a small change in parameter space. In contrast, small-$\tau$ downstream initializations require large parameter movement to fit the data in the delayed rich regime.
@@
Upstream. When $\delta_{k}\gg 0$, $M_{k}\approx\delta_{k}\mathbf{I}_{d}$, and the dynamics are approximately $\partial_{t}\hat{\beta}_{k}=-\delta_{k}\mu_{k}^{-1}(\mathbf{I}_{d}-\hat{\beta}_{k}\hat{\beta}_{k}^{\intercal})\xi_{k}$ and $\partial_{t}\mu_{k}=-\delta_{k}\hat{\beta}_{k}^{\intercal}\xi_{k}$. Again, both the direction and magnitude of $\beta_{k}$ change. However, unlike the balanced setting, in this setting $M_{k}$ is independent of $\beta_{k}$ and stays constant through training. Yet, as $\beta_{k}$ change in direction, so can $C$, and thus the NTK. This setting is unique, because it is rich due to a changing activation pattern, but the dynamics do not move far in parameter space. Furthermore, unlike in the balanced scenario where scale adjusts the speed of radial dynamics, here it regulates the speed of directional dynamics, with vanishing initializations prompting an extremely fast alignment phase, as observed in Fig. 1 and in Fig. 5.
@@
Figure 6: Impact of upstream initializations in practice. Here we provide evidence that an upstream initialization (a) drives feature learning through changing activation patterns, (b) promotes interpretability of early layers in CNNs, (c) reduces the sample complexity of learning hierarchical data, and (d) decreases the time to grokking in modular arithmetic. In these experiments, we regulate the first layer’s learning speed relative to the rest of the network by dividing its initialization by $\alpha$. For models without normalization layers, we also scale the last layer’s initialization by $\alpha$ to preserve the input-output map. $\alpha=1$ represents standard parameterization, while $\alpha\gg 1$ and $\alpha\ll 1$ correspond to upstream and downstream initializations, respectively. See Section D.3 for details.
Connections to infinite-width. Our study of learning regimes in finite-width two-layer ReLU networks as a function of the overall and relative scale is consistent with existing infinite-width analysis of feature learning. For example, in Luo et al. [17] they consider a network $f(x)=\frac{1}{\alpha}\sum_{k=1}^{h}a_{k}\sigma(w_{k}^{\intercal}x)$ with weights initialized as $a_{k}\sim\mathcal{N}(0,\beta_{a}^{2})$ and $w_{k}\sim\mathcal{N}(0,\beta_{W}^{2}\mathbf{I}_{d})$ as width $h\to\infty$. They obtain a phase diagram at infinite width capturing the dependence of learning regime on the overall function scale $\beta_{a}\beta_{W}/\alpha$ and the relative initialization scale $\beta_{a}/\beta_{W}$, each suitably normalized as a function of width. The resulting phase portrait is analogous to ours in Fig. 1 (b), where we use the conserved quantity $\delta$ rather than the relative scale $\beta_{a}/\beta_{W}$. Specifically, there is a lazy regime that includes the NTK parameterization, which is always achieved at large scale (as in the large-$\tau$ regions of Fig. 1 (b)), but is also achieved at small scale if the first layer variance is sufficiently larger than the second (as in the downstream initializations at small $\tau$ in Fig. 1 (b)). On the other side of the phase boundary is the infinite-width analog of rapid rich learning, where all neurons condense to a few directions. This is induced either at small function scale, or at larger scales if $\beta_{a}/\beta_{W}$ is sufficiently large, such that $W$ learns fast enough relative to $a$. The phase boundary, in turn, which exists only at infinite width, contains a range of parametrizations, including the mean-field parametrization. More broadly, across width-dependent parametrizations, the random initialization of weights induces a distribution over per-neuron conserved quantities. While the distinction between the NTK and the mean-field parametrizations has been extensively studied, both lead to the same distribution of per-neuron conserved quantities, which is zero in expectation with a non-vanishing variance. A more thorough study of what role the distribution of per-neuron conserved quantities plays in feature learning at finite-widths is left to future work.